# TransferLight: Zero-Shot Traffic Signal Control on any Road-Network

### v1

@@ 1 Introduction
Another significant issue in current RL-based approaches is that the variability of traffic dynamics is not adequately accounted for during the training process. The majority of methods employ identical spatio-temporal traffic patterns across all training episodes (Wei et al. 2018, 2019a). While these models may exhibit impressive performance within this constrained setting, they typically suffer from substantial performance degradation when confronted with real-world variability (Zheng et al. 2019; Yoon et al. 2021). This performance decline can be attributed to overfitting and the drastically constrained exploration space during training (Jiang, Kolter, and Raileanu 2024). The limited exposure to diverse traffic scenarios during the learning phase results in models that lack robustness and adaptability to the complex and dynamic nature of real-world traffic conditions (Korecki, Dailisan, and Helbing 2023). Consequently, these models struggle to generalize effectively to the multifaceted and often unpredictable traffic patterns encountered in practical applications, highlighting a critical gap between laboratory performance and real-world efficacy.
@@
The limitations of existing traffic signal control (TSC) approaches, particularly their inability to generalize across intersection configurations and traffic conditions, necessitate a more robust and flexible solution. [^1] The ideal solution would be a model capable of maintaining consistent, high-quality performance across a wide spectrum of road-networks and traffic dynamics. Once the general control policy is obtained, it can be applied to any (urban) environment. We present TransferLight, a novel model that addresses these challenges by leveraging graph-structured representations and advanced training techniques. Our contributions include
@@ 2 Priliminaries
In fully cooperative Markov Games, the global reward is equivalent to individual rewards ($R^{t}=R_{v}^{t},\forall v\in\mathcal{V}$) or a team average ($R^{t}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}R_{v}^{t}$). While the former entails aligned goals for individual agents, the latter allows agents to pursue distinct objectives that contribute to the overall team benefit. Since individual and global rewards are functions of joint actions, value functions also depend on the joint policy $\pi=\{\pi_{v}\mid v\in\mathcal{V}\}$ of all agents. Based on the reward definition, we can define an individual state-value function $\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R_{v}^{t+k}\mid\mathcal{C}^{t}]$ or a global state-value function $\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R^{t+k}\mid\mathcal{C}^{t}]$. The objective is to find an optimal joint policy $\pi^{\ast}$ that maximizes the expected discounted sum of global rewards:
@@ 4 Graph-Structured State Encoding
While often overlooked in previous studies (Zheng et al. 2019; Oroojlooy et al. 2020; Zang et al. 2020; Yoon et al. 2021), the length of lanes or segments plays a crucial role in traffic dynamics. Our approach employs a uniform and constant segment length $ds$, thereby streamlining the input feature set compared to related works (Devailly, Larocque, and Charlin 2022, 2024). This design choice allows the policy to implicitly learn length-related characteristics, including segment capacity, enhancing the model’s adaptability to diverse road networks compared to prior work (Wei et al. 2019a). However, $ds$ has to be small enough [^8] Note that $\frac{\ell}{ds}$ can have a remainder, which we drop. The impact is minor for any reasonable choices of $ds$, as the cut-off is done at the end of the lane (maximally distant to the intersection). We choose $ds$ to be $10$ meters for our experiments. to minimise the impact of local degeneracies, as discussed in Section 3.
@@
To break the permutation invariance of the segment set, we define the centre of the intersection as a reference point and induce a positional encoding on the segments relative to that point. We use a one-dimensional sinusoidal positional encoding (Vaswani et al. 2017) along segments on each lane and over lanes. Instead of additive fusion (Vaswani et al. 2017), we concatenate the positional information with the density of the segment. This preserves both identities, which improves expressiveness without the need of separate processing (Yu et al. 2023). Thus, we can define the segment feature vector by $\mathbf{h}_{s}=[\rho(s)\|\operatorname{pe}(s)\|\bar{\rho}_{\ell}]^{\top}\in\mathbb{R}^{d}$ be the feature vector of a segment $s\in\ell$.
@@
where $\mathbf{a}_{m\phi},\mathbf{b}_{m}\in\mathbb{R}^{d^{\prime}}$ and $\mathbf{W}_{m},\tilde{\mathbf{W}}_{m}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ are learnable weights. Attention weights are computed as in Eq. 8 but normalised over $\mathcal{M}_{v}$ instead. Contrary to Veličković et al. (2017), we embed node and edge features separately, which reduces the model complexity while still preserving expressiveness. Furthermore, we use the edge features and the initial phase flag only to compute the attention scores. Hence, the model can use $\gamma$ to weight movement features during aggregation, but they do not infer otherwise with the movement information. As we use a directed acyclic graph, we do not face the identity issue discussed in general edge-based graph attention (Wang, Chen, and Chen 2021). This form of aggregation also preserves permutation invariance. In contrast to the level before, this is an important property for the encoding of phases, as they should be orientation independent (Zheng et al. 2019). The obtained phase node representations are further leveraged in an intra-level propagation phase, as discussed next.
@@ 6 Experiments
The primary objective of our experiments is to show the ability of TransferLight to transfer its control policy to novel scenarios without requiring any kind of re-training or fine-tuning. In all experiments, TransferLight is trained on randomly generated road-networks with random traffic dynamics and tested on a yet unseen benchmark. This allows us to quantify the generalisability of our method explicitly. In Section 6.1, we analysed various performance measures on multiple benchmarks (test scenarios) with several well-known baselines. As arterial scenarios are of specific interest for the community (Wei et al. 2019a), we conduced a detailed investigation of our model’s generalisability on such scenario types (see Section 6.2). The software specifications of our implementations can be found in our open-sourced code.
@@ 6.2 Arterial Signal Progression
Figure 5: Average Travel Time on Cologne8 over 3600 simulated time steps.
@@ 7 Conclusion
We presented a novel framework designed for robust generalization across road-networks, diverse traffic conditions and intersection geometries. Our method can scale to any road-network through a decentralized multi-agent approach with global rewards and state transition priors to ensure proactive decisions. We used a hierarchical, heterogeneous, and directed graph neural network to encode any intersection geometry, which we train using a novel log-distance reward function. Generalization is further fostered by domain randomization during training. Through domain randomization during training, we additionally enhance generalization capabilities. This is particularly valuable for real-world applications where traffic conditions can vary significantly due to events, road closures, or long-term changes in urban mobility patterns.
@@ A.1 Implementation Details
A.2 Further Experiments
@@
Limits of Generalisability
The ability to generalise is of course facing limits at some range of problem complexity. We performed an additional experiment on ingolstadt21 comprising $21$ intersections in a narrow urban environment. Figure 10 compares our method to various baselines under different performance measures on this benchmark scenario. After around 1200 time steps, TransferLight with either head starts diverging into a suboptimal sequence of phases. On the long run, this leads to congestions, which in turn lead to performance decreases among all measures. We dedicate our future work to prevent such situations to occur (under reasonable traffic demands).

### v2

@@ 1 Introduction
Another significant limitation in current RL-based approaches is the insufficient consideration of traffic dynamics’ variability during training. The majority of methods employ identical spatio-temporal traffic patterns across all training episodes (Wei et al. 2018, 2019a). While these models may exhibit impressive performance within this constrained setting, they typically suffer from substantial performance degradation when confronted with real-world variability (Zheng et al. 2019; Yoon et al. 2021). This performance decline can be attributed to overfitting and the drastically constrained exploration space during training (Jiang, Kolter, and Raileanu 2024). The limited exposure to diverse traffic scenarios during the learning phase results in models that lack robustness and adaptability to the complex and dynamic nature of real-world traffic conditions (Korecki, Dailisan, and Helbing 2023). Consequently, these models struggle to generalize effectively to the multifaceted and often unpredictable traffic patterns encountered in practical applications, highlighting a critical gap between laboratory performance and real-world efficacy.
@@
The limitations of existing traffic signal control (TSC) approaches, particularly their inability to generalize across intersection and traffic conditions, necessitate a more robust and flexible solution. [^1] The ideal solution would be a model capable of maintaining consistent, high-quality performance across a wide spectrum of road-networks and traffic dynamics. Once the general control policy is obtained, it can be applied to any (urban) environment. We present TransferLight, a novel model that addresses these challenges by leveraging graph-structured representations and advanced training techniques. Our contributions include
@@ 2 Priliminaries
In fully cooperative Markov Games, the global reward is equivalent to individual rewards ($R^{t}=R_{v}^{t},\forall v\in\mathcal{V}$) or a team average ($R^{t}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}R_{v}^{t}$). While the former entails aligned goals for individual agents, the latter allows agents to pursue distinct objectives that contribute to the overall team benefit. The objective is then to find a joint policy $\pi=\{\pi_{v}\mid v\in\mathcal{V}\}$ that maximizes the expected discounted sum of global future rewards:
@@ 4 Graph-Structured State Encoding
While often overlooked in previous studies (Zheng et al. 2019; Oroojlooy et al. 2020; Zang et al. 2020; Yoon et al. 2021), the length of lanes or segments plays a crucial role in traffic dynamics. Our approach employs a uniform and constant segment length $ds$, thereby streamlining the input feature set compared to related works (Devailly, Larocque, and Charlin 2022, 2024). This design choice allows the policy to implicitly learn length-related characteristics, including segment capacity, enhancing the model’s adaptability to diverse road networks compared to prior work (Wei et al. 2019a). However, $ds$ has to be small enough to minimise the impact of local degeneracies, as discussed in Section 3.
@@
To break the permutation invariance of the segment set, we define the centre of the intersection as a reference point and induce a positional encoding on the segments relative to that point. We use sinusoidal positional encoding (Vaswani et al. 2017) along segments on each lane and over lanes. Instead of additive fusion (Vaswani et al. 2017), we concatenate the positional information with the density of the segment. This preserves both identities, which improves expressiveness without the need of separate processing (Yu et al. 2023). Thus, we can define the segment feature vector by $\mathbf{h}_{s}=[\rho(s)\|\operatorname{pe}(s)\|\bar{\rho}_{\ell}]^{\top}\in\mathbb{R}^{d}$ be the feature vector of a segment $s\in\ell$.
@@
where $\mathbf{a}_{m\phi},\mathbf{b}_{m}\in\mathbb{R}^{d^{\prime}}$ and $\mathbf{W}_{m},\tilde{\mathbf{W}}_{m}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ are learnable weights. Attention weights are computed as in Eq. 8 but normalised over $\mathcal{M}_{v}$ instead. Contrary to Veličković et al. (2017), we embed node and edge features separately, which reduces the model complexity while still preserving expressiveness. Furthermore, we use the edge features and the initial phase flag only to compute the attention scores. Hence, the model can use $\gamma$ to weight movement features during aggregation, but it does not make any further inferences from the movement information. As we use a directed acyclic graph, we do not face the identity issue discussed in general edge-based graph attention (Wang, Chen, and Chen 2021). This form of aggregation also preserves permutation invariance. In contrast to the level before, this is an important property for the encoding of phases, as they should be orientation independent (Zheng et al. 2019). The obtained phase node representations are further leveraged in an intra-level propagation phase, as discussed next.
@@ 6 Experiments
The primary objective of our experiments is to show the ability of TransferLight to transfer its control policy to novel scenarios without requiring any kind of re-training or fine-tuning. In all experiments, TransferLight is trained on randomly generated road-networks with random traffic dynamics and tested on a yet unseen benchmark. This allows us to quantify the generalisability of our method explicitly. In Section 6.1, we analysed various performance measures on multiple benchmarks (test scenarios) with several well-known baselines. As arterial scenarios are of specific interest for the community (Wei et al. 2019a), we conduced a detailed investigation of our model’s generalisability on such scenario types (see Section 6.2). The software specifications of our implementations can be found in our open-sourced code [^11] https://github.com/johSchm/TransferLight.
@@ 6.2 Arterial Signal Progression
Figure 5: Average Travel Time on Cologne3 over 3600 simulated time steps.
@@ 7 Conclusion
We presented a novel framework designed for robust generalization across road-networks, diverse traffic conditions and intersection geometries. Our method can scale to any road-network through a decentralized multi-agent approach with global rewards and state transition priors to ensure proactive decisions. We used a heterogeneous and directed graph neural network to encode any intersection geometry, which we train using a novel log-distance reward function. Generalization is further fostered by domain randomization during training. This is particularly valuable for real-world applications where traffic conditions can vary significantly due to events, road closures, or long-term changes in urban mobility patterns.
@@ A.1 Implementation Details
For our random training environments, we simulated passenger cars as vehicles only. The same is true for cologne1, cologne3 and cologne8. In addition to vanilla passenger cars, ingolstadt1, ingolstadt7 and ingolstadt21 also include buses.
Segment Length
We choose a segment length $ds$ of $10$ meters for our experiments. In theory, $ds$ is only upper bounded to $\ell$. However, we argue in favour of tighter bounds in practice. Using segment densities as inputs factors out the number of vehicles and enables the encoder to learn segment lengths implicitly (if needed). This requires a globally consistent segment length $ds$ (as mentioned in Section 4). Hence, we drop possible remainders of $\frac{\ell}{ds}$. The impact is minor for any reasonable choices of $ds$, as the cut-off is done at the end of the lane (maximally distant to the intersection). This approach introduces a technical upper bound to the choice of $ds$, as short lanes with $\ell<ds$ would be dropped. Furthermore, choosing $ds$ smaller than the average vehicle length $h$ does not offer more valuable information for the encoding. Therefore, we have $h\leq ds\leq\min_{\ell\in\mathcal{I}\cup\mathcal{O}}\ell$.
@@
A.2 Theory
Runtime Complexity
Signals on the segment level are embedded in parallel and aggregated by Eq. 9. Let $S_{v}:=\sum_{\ell\in\mathcal{I}_{v}\cup\mathcal{O}_{v}}\frac{\ell}{ds}$ be the number of segments at $v$. Due to the quadratic complexity of the attention mechanism in each of the three encoding layers, we get
$\mathcal{O}\left(S_{v}|\mathcal{M}_{v}|+|\mathcal{M}_{v}||\Phi_{v}|+\Phi_{v}^{2}+P_{v}\right),$ (14)
where $P_{v}:=\sum_{\ell\in\mathcal{I}_{v}\cup\mathcal{O}_{v}}|\mathcal{M}_{\rightarrow\ell}|+|\mathcal{M}_{\leftarrow\ell}|$ is the number of movements required to estimate the transition priors in Eq. 7. Note that $P_{v}\ll|\mathcal{M}|$ as only the prior only considers movements connecting two intersections. For a reasonable segment length $ds$, $S_{v}|\mathcal{M}_{v}|$ is the dominant term in Eq. 14. As each intersection is controlled by weight-tied agents, policies can be computed in parallel for all $v$, disentangling the number of intersections from the runtime complexity.
A.3 Ablation Studies
@@
Refer to caption $\operatorname{pe}$ $\gamma_{m\phi}$ $J_{\phi\phi^{\prime}}$ $\bar{\rho}_{\ell}$ ATT ($\downarrow$)
$92.13$$\pm 58.18$
$144.52$$\pm 107.83$
$92.15$$\pm 58.22$
$92.80$$\pm 58.23$
$\mathbf{91.71}$$\pm\mathbf{56.51}$
Figure 10: Ablation study of positional encoding $\operatorname{pe}$, edge features $\gamma_{m\phi}$ and $J_{\phi\phi^{\prime}}$, and our transition prior $\bar{\rho}_{\ell}$ comparing Average Travel Time (ATT) on cologne8. Left: visualization of the cologne8 network. Right: ablation results.
Impact of Positional Encoding
When lane segments are relatively short on average, the inclusion of positional encoding at the segment level has limited influence on model performance. This condition is particularly prevalent in urban environments, where signalized intersections are typically located in close proximity to one another. Under such circumstances, positional encoding can be effectively disregarded. Evidence supporting this claim is presented in Fig. 10, where the omission of positional encoding resulted in only a negligible increase in average travel time.
Impact of Transition Prior
The removal of the state transition prior during both training and testing resulted in the second-largest performance degradation observed in our ablation study (see Fig. 10). The transition prior plays a pivotal role in enabling the model to make proactive decisions rather than merely reactive ones, a capability that is especially critical in dense, urban environments. However, the observed impact of this prior was less substantial than initially anticipated. This discrepancy warrants further investigation, and we plan to explore the underlying reasons in greater detail in future work.
Impact of Edge Features
To introduce a measure of similarity between traffic signal phases (in terms of their green signal overlap), we employed the Jaccard index at the phase-to-phase level. As demonstrated in Fig. 10, this feature provided only a marginal improvement in performance. At the movement-to-phase level, we incorporated edge feature encodings to represent whether a movement was prohibited, protected, or permitted. By a wide margin, this feature had the most significant impact on overall performance. We hypothesize that this added inductive bias enables the model to cluster movements more effectively before decoding them into phase energy distributions. Consequently, the model focuses primarily on traffic densities that can be directly alleviated by selecting specific phases. Investigating this claim more thoroughly is another key point for potential follow-up work.
@@
A.4 Limits of Generalisability
The ability to generalise is of course facing limits at some range of problem complexity. We performed an additional experiment on ingolstadt21 comprising $21$ intersections in a narrow urban environment. Figure 9 compares our method to various baselines under different performance measures on this benchmark scenario. After around 1200 time steps, TransferLight with either head starts diverging into a suboptimal sequence of phases. On the long run, this leads to congestions, which in turn lead to performance decreases among all measures. We dedicate our future work to indicate the causing factors and prevent such situations to occur (under reasonable traffic demands).